OpenStudy (anonymous):

I need to differentiate this equation using derivative rules: sqrt3x x^2

OpenStudy (anonymous):

\[\sqrt{3x}x{2}\]

zepdrix (zepdrix):

\[\large \left(\sqrt{3x}\cdot x^2\right)' \qquad=\qquad \color{royalblue}{\left(\sqrt{3x}\right)'}x^2+\sqrt{3x}\color{royalblue}{\left(x^2\right)'}\] We have the product of two things involving x, so we start by applying the product rule. The blue terms are the ones we still need to differentiate, that's what the prime notation is telling us. Understand the setup so far?

OpenStudy (anonymous):

yup

zepdrix (zepdrix):

Since the sqrt3 is a constant, we don't need to worry about it in the differentiation process. \[\large \left(\sqrt{3x}\right)' \qquad=\qquad \sqrt3\left(\sqrt x\right)'\] If you don't have this derivative memorized, you can change the term to a rational expression using exponent rules. \[\large \sqrt3\left(\sqrt x\right)'\qquad=\qquad \sqrt{3}\left(x^{1/2}\right)'\]From here, you would apply the power rule to the x term. Confused about any of this? :o

zepdrix (zepdrix):

Oh maybe we shouldn't do that.. separating the 3 and x... Because it's a good idea to get some practice with the chain rule.. hmm my bad

OpenStudy (anonymous):

well we haven't learned the chain rule yet so its ok

zepdrix (zepdrix):

Oh ok :)

zepdrix (zepdrix):

Do you know the derivative of \(\large x^{1/2}\) ?

OpenStudy (anonymous):

no i don't.

OpenStudy (anonymous):

also how did you separate sqrt 3?

OpenStudy (anonymous):

i know that x2 = 2x. by using the power rule

zepdrix (zepdrix):

\[\large \sqrt{3x}\qquad=\qquad \sqrt3\cdot\sqrt x\] Since sqrt3 is a constant, we're allowed to pull it outside of the differentiation process, so I took it outside of those brackets.

zepdrix (zepdrix):

\[\large \left(\sqrt{3x}\cdot x^2\right)' \qquad=\qquad \color{royalblue}{\left(\sqrt{3x}\right)'}x^2+\sqrt{3x}\color{orangered}{\left(2x\right)}\]Ok very good :) That takes care of that one.

OpenStudy (anonymous):

so would x^1/2 equal to... 1/2x^1/4?

OpenStudy (anonymous):

oh right. duh the constant rule. ok

zepdrix (zepdrix):

\[\large x^{1/2}\] Using the `Power Rule for Derivatives`: Step 1, we'll bring the power down as a factor in front. Step 2, we'll subtract 1 from the exponent.\[\large \frac{1}{2}x^{1/2-1}\]I think you did your subtraction wrong or something there.

OpenStudy (anonymous):

yeah I did lol.

OpenStudy (anonymous):

so now i have to take the derivative of sqrt3x right?

zepdrix (zepdrix):

Yes but we were pulling the sqrt3 outside to make our derivative easier.

zepdrix (zepdrix):

\[\large \left(\sqrt{3x}\cdot x^2\right)' \qquad=\qquad \sqrt{3}\color{royalblue}{\left(\sqrt{x}\right)'}x^2+\sqrt{3x}\color{orangered}{\left(2x\right)}\]Just need to deal with the blue part.

OpenStudy (anonymous):

right

zepdrix (zepdrix):

Did you get that exponent figured out correctly? It's not 1/4 :)

OpenStudy (anonymous):

yeah i did but why did you ask?

OpenStudy (anonymous):

I don't remember where x^1/2 was in the question...

zepdrix (zepdrix):

\[\large \left(\sqrt{3x}\cdot x^2\right)' \qquad=\qquad \sqrt{3}\color{royalblue}{\left(x^{1/2}\right)'}x^2+\sqrt{3x}\color{orangered}{\left(2x\right)}\]

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

ok but now wouldn't it be just \[x ^{-1/2}\]

OpenStudy (anonymous):

when you do 0.5-1

zepdrix (zepdrix):

Yes good :) With a 1/2 in front.

zepdrix (zepdrix):

\[\large \left(\sqrt{3x}\cdot x^2\right)' \qquad=\qquad \sqrt{3}\color{orangered}{\left(\frac{1}{2}x^{-1/2}\right)}x^2+\sqrt{3x}\color{orangered}{\left(2x\right)}\]

zepdrix (zepdrix):

And that is our derivative! Yay team! From there, your teacher might expect you to clean it up a little bit. But whatev.

OpenStudy (anonymous):

that's it? it doesn't simplify anymore?

zepdrix (zepdrix):

\[\large x^{-1/2}\qquad=\qquad \frac{1}{x^{1/2}}\qquad=\qquad \frac{1}{\sqrt x}\] So pulling the 1/2 out of the first term, we could rewrite it as,\[\large \left(\sqrt{3x}\cdot x^2\right)' \qquad=\qquad \frac{\sqrt{3}}{2}\color{orangered}{\left(\frac{1}{\sqrt x}\right)}x^2+\sqrt{3x}\color{orangered}{\left(2x\right)}\]And then ... ya you could do some simplification beyond that maybe :p

OpenStudy (anonymous):

why did the 2 go under sqrt 3

zepdrix (zepdrix):

\[\large \sqrt3 \cdot\frac{1}{2}\qquad=\qquad \frac{\sqrt3}{1}\cdot \frac{1}{2}\qquad=\qquad \frac{\sqrt3}{2}\]

OpenStudy (anonymous):

oh yeah forgot about that!

OpenStudy (anonymous):

ok thanks so much!